W: “Intertwined” and “ahead of his time”

Doron ZEILBERGER^{1}

It is hard for me to write something solicited about W. He was the better half of “WZ pair” and “WZ theory”. If I had to choose one word to describe him, it would be “intertwined”. His beloved wife of sixty years, Ruth Tumen W., said in her eulogy, “for him research and teaching was intertwined”. I wholeheartedly agree, but not only teaching and research. The computational and the conceptual, the pure and applied, the big picture and the nitty-gritty details of Fortran or Mathematica code, research and exposition (and boy was he a great expositor!) the professional and the personal, and so on and so forth.

He was also way ahead of his time! When computers were just starting to get used, he edited, in the lated fifties, together with Anthony Ralston, a seminal “cookbook” telling people how to solve mathematical problems with computers, complete with flowcharts and pseudocode (before Fortran!). Fifteen years later, he cowrote, with Albert Nijenhuis, the bible of combinatorial algorithms, that intertwined theory and implementation. When the internet started, he pioneered the Electronic J. of Combinatorics, a free(!) on-line journal. He convinced the publishers of his numerous books to let them post his books freely on the internet, thereby being a pioneer of open-access, and open source-code. When Tim Gowers asked him to write the entry on Enumerative and Algebraic combinatorics he politely refused, but suggested that he write instead an entry on “Experimental Mathematics”, thereby introducing this avant-garde subject into the canon of mathematics.

And his mathematics was divine! Please watch my lecture:

http://youtu.be/rnFWmswEFFA (part 1), http://youtu.be/8NwFQGO3tzU (part 2) .

1 Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen

Rd., Piscataway, NJ 08854-8019, USA. zeilberg at math dot rutgers dot edu , http://www.math.rutgers.edu/~zeilberg/ . May 22, 2012.